OpenStudy (anonymous):

The area of a rectangle is 102 square feet and the perimeter is 46 feet. Find the dimensions of the rectangle

OpenStudy (zzr0ck3r):

 The are is given by \(A=x*y\) and perimeter is \(P=2x+2y\) So we have \(102=xy\implies x=\frac{102}{y}\) And we also have \(2x+2y=46\) So we plug \(x=\frac{102}{y}\) for \(x\) and solve for \(y\). \(2\frac{102}{y}+2y=46\implies \frac{204}{y}+2y=46\implies 204+2y^2=46y\) So we have \(2y^2-46y+204=0\). Can you solve this equation for \(y\)?